OpenStudy (daniellelovee):

The equation of an ellipse is shown. (x-43)^2/225 + (y-68)^2/169 =1 What is the length of its major axis? A. 15 B. 26 C. 30 D. 43

OpenStudy (xapproachesinfinity):

first rewrite it as \[\frac{(x-43)^2}{15^2}+\frac{(y-68)^2}{13^2}=1\] so the major axis is 2a=2x15=30